# A Homothetic Gauge Theory

### v1

@@ 7 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
Given a closed smooth $(d-1)$-dimensional hypersurface $S$ in ${\mathbb{R}}^{d}$, we denote by $\delta_{n}$ a sequence of regularizations of the Dirac delta function $\delta_{S}$ (not to be confused with the codifferential in the previous sections), such that $\lim_{n\to\infty}\delta_{n}=\delta_{S}$. Then, in the generalized Laplace equation (49), we choose (see Appendix B for justification):
@@
We observe that, in the singular limit $n\to\infty$, the regularization $\delta_{n}$ approaches the Dirac distribution $\delta_{S}$, supported on a closed hypersurface $S\subset\mathbb{R}^{d}$. In this limit, the terms involving derivatives of $\delta_{n}$ become highly localized near $S$, and act to impose both Dirichlet and Neumann conditions simultaneously on the surface.
As a result, equation (56) defines two decoupled Cauchy problems in $\mathbb{R}^{d}$: one on the interior and one on the exterior of $S$. Unless one solution is the harmonic continuation of the other, a discontinuity (in the function or its gradient) appears at $S$. If only a Dirichlet condition $\phi_{S}=\phi_{d}$ is imposed, a discontinuity in the normal derivative $\nabla\phi_{S}\cdot\mathbf{n}$ may appear across $S$. Conversely, if only the Neumann condition $\nabla\phi_{S}\cdot\mathbf{n}=\nabla\phi_{d}\cdot\mathbf{n}$ is enforced, a jump in the potential $\phi_{S}$ itself can occur. These phenomena are closely related to classical potential theory, particularly the behavior of single-layer and double-layer potentials [20, 21].
Specifically, a single-layer potential yields a continuous potential across the interface but typically introduces a discontinuity in its normal derivative. In contrast, a double-layer potential maintains continuity of the normal derivative but allows for a jump in the potential itself [5]. The behavior of equation (56) in the limit $n\to\infty$ is thus analogous to a combined layer potential, where both Dirichlet and Neumann data are enforced, resembling classical integral formulations for boundary value problems in potential theory.
Interestingly, the structure of equation (56) also permits a modular strategy for solving classical boundary value problems. Suppose one intends to solve a pure Dirichlet problem—either on the interior or exterior of $S$. In this case, one may omit the gradient-coupling term and instead consider the equation:
@@
Once $\phi_{n}$ is obtained from equation (57), one automatically gains access to the normal derivative $\nabla\phi_{n}\cdot\mathbf{n}$ on $S$. In the context of elliptic boundary value problems, this is equivalent to applying the Dirichlet-to-Neumann (DtN) map (not to be confused with the homothetic transformation matrix in the previous sections), often denoted by
$\Lambda:\phi_{n}|_{S}\mapsto\nabla\phi_{n}\cdot\mathbf{n}|_{S}.$
Now, suppose we set $\nabla\phi_{d}:=\nabla\phi_{n}$; then, the term $(\nabla\phi_{n}-\nabla\phi_{d})$ in the original equation (56) vanishes identically, and the full equation is satisfied. Thus, one has in effect solved the complete model under a self-consistent assumption that aligns the constructed $\phi_{d}$ with the solution gradient on $S$.
A similar argument holds for solving Neumann problems. In this case, one may omit the potential penalty term and retain only the gradient-coupling component:
@@
This imposes a Neumann-type boundary condition by penalizing discrepancies between the solution gradient and the prescribed gradient field $\nabla\phi_{d}$. Upon obtaining a solution $\phi_{n}$, one can define $\phi_{d}:=\phi_{n}$ on $S$, and again observe that the residual term $(\phi_{n}-\phi_{d})$ in equation (56) vanishes. Hence, the full model is satisfied under the assumption of gradient-data consistency.
These selective reductions of the model not only simplify numerical treatment but also clarify the role of the original PDE as a unified interface formulation capable of encoding a range of classical boundary conditions through regularized volume penalties. The correspondence with the DtN operator also invites further interpretation of the interface as a generalized boundary operator in the sense of nonlocal elliptic theory [6, 25].
8 A Non-Singular Model for a Point Electric Charge
@@
diverges due to the integrable singularity at the origin. This divergence has been a foundational issue in classical field theory, prompting numerous attempts at regularization, ranging from distributional reformulations [11, 4] to the introduction of extended charge models [7, 2].
In the framework developed in the previous section, a new perspective emerges. We propose a regularized, non-singular model for a point electric charge using the generalized Laplace equation with Dirichlet-type penalization:
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0,$ (57 revisited)
where $\delta_{n}$ is a regularization of the Dirac delta function supported on the surface of a ball, $\delta_{\partial B(R)}$, and $\phi_{d}$ is the prescribed Dirichlet potential.
We consider a spherical interface $\partial B(R)\subset\mathbb{R}^{d}$, and solve the generalized equation (57) in the exterior domain $\mathbb{R}^{d}\setminus B(R)$. We set the Dirichlet data on the boundary to be
$\phi_{d}=\frac{C}{R},$
enforcing a Coulomb-like condition at $r=R$. Due to the highly localized nature of the penalty term, the solution $\phi_{n}$ adjusts sharply near the interface to match this boundary value, and decays as $\phi_{n}(r)\sim\frac{C}{r}$ for large $r$, consistent with the behavior of the classical potential in the far field.
$2R$$\dfrac{C}{R}=\phi_{d}$$\phi_{n}(r)\sim\dfrac{C}{r}$nterior solution doesn’taffect exterior solution
Figure 4: A non-singular point charge model with a spherical interface at $r=R$. The interior potential is decoupled from the exterior, where the solution decays as $\phi_{n}\sim\frac{C}{r}$.
Importantly, the structure of the equation ensures a natural decoupling of the interior and exterior domains. In the limit $n\to\infty$, the penalization enforces the boundary condition only on $\partial B(R)$, and the interior and exterior solutions become independent. Thus, the behavior of the solution in $\mathbb{R}^{d}\setminus B(R)$ is determined entirely by the enforced Dirichlet value and is unaffected by the field inside the ball.
We now can safely consider the limiting case $R\to 0$, by which, the ball $B(R)$ approaches a point, and the prescribed boundary potential $\phi_{d}=C/R$ diverges, mimicking the classical singular behavior at the origin. However, this divergence is confined to an infinitesimally small region surrounding the origin, and—crucially—does not contribute to the total field energy. Since the equation is solved only in the exterior domain, and the potential inside the vanishing ball plays no role in the exterior solution, the infinite boundary value introduces no singularity into the energy functional. The energy,
[…]
remains finite in the limit $R\to 0$, as the singular core is effectively excluded from the domain of integration.
This construction achieves what the classical model could not: a physically faithful, asymptotically Coulombic potential associated with a point-like source, but without the divergence of vacuum energy. The source is interpreted not as a singular point but as the limit of a penalized interface of vanishing size, carrying an enforced boundary potential. The infinite energy problem is thereby resolved through a combination of geometrically localized penalization and the decoupling of interior and exterior fields.
@@ 9 Conclusions
This work has introduced a novel gauge theory of scale, placing local homotheties on an equal footing with the familiar internal symmetries that underpin the Standard Model. By constructing a formalism where local scale transformations act linearly, we have developed a self-contained and consistent framework with significant physical implications. The principal results of this investigation demonstrate the theory’s viability and its potential to resolve long-standing issues in classical physics:
  • A generalized Weyl-integrable geometry emerges naturally from our framework, where the spacetime metric dynamically responds to a scalar gauge field, $\lambda$, that dictates the local scale.
  • We established a homothetic de Rham complex and an associated Hodge decomposition. This mathematical backbone guarantees that the resulting field equations are well-posed and that their physical solutions are unique.
  • Applying the theory to electromagnetism yields homothetically covariant sourceless Maxwell equations. These equations have some additional terms coupling the electromagnetic field with the gradient of the scale gauge field.
  • A key achievement is a resolution to the century-old problem of the infinite self-energy of a point charge. Our model replaces the singularity with a regularized, penalized boundary condition on an infinitesimal interface—a result that emerges directly from the theory’s geometric foundations, not from ad-hoc modifications.
  • The entire gauge structure was shown to have a rigorous interpretation within the established framework of Cartan geometry, clarifying its deep connection to the conformal group and providing a basis for further geometric analysis.
Taken together, these findings present homothetic gauge theory as more than a mathematical curiosity; it is a workable extension of classical field theory. The future avenues for exploration are rich and varied. Integrating HGT with metric dynamics could offer a novel path toward modifying general relativity, where cosmological phenomena might be linked to gauge artifacts of the scale field. Furthermore, coupling the homothetic connection to non-Abelian gauge groups could unveil new scale-dependent physics relevant to the strong and weak interactions. The theoretical and computational tools developed here provide a solid foundation for pursuing these exciting possibilities.
@@ A.1 Conformal Cartan Geometry
  • the doubled bundle $TM\oplus TM$ of Section 2 is the direct sum of two tractor bundles carrying weights 00 and $-1$, respectively.
Hence all operators in §$\S$ 2 are functorial in Cartan geometry.
@@ A.3 Interpretation of the Fixed Vector $X_{d}$
The canonical embedding $X\mapsto\bar{X}=\bigl{(}X~{}~{}X_{d}\bigr{)}^{\!\top}$ used in the main text corresponds, in the Cartan picture, to the choice of a weight $-1$ compensator tractor $X_{d}$. Its presence linearizes local scale transformations, exactly as a compensator field does in Weyl gravity.
@@ Appendix B Finding $\lambda$ for imposition of boundary conditions
Here we find the $\lambda(x)$ in a way that allows us to apply boundary conditions. The method presented here is not a complete solution to the mathematical problem of finding $\lambda$, but the $\lambda$ found is acceptable for sufficiently regular boundaries.
@@
[…]
@@ Appendix C Principal $\mathcal{H}$-bundle formulation and curvature
Let $\mathcal{H}\cong({\mathbb{R}}_{>0},\cdot)$ be the homothetic group acting on each doubled fibre by block-diagonal multiplication. [^3] Locally—on a chart $(U,x)$—a homothety is specified by a smooth function $\lambda\in C^{\infty}(U)$, the action being $v\mapsto e^{\lambda(x)}v$. Pointwise multiplication turns $\mathcal{H}$ into a $1$-dimensional abelian Lie group with Lie algebra $\mathfrak{h}\cong{\mathbb{R}}$ under addition. Choose a good open cover $\{U_{\alpha}\}_{\alpha\in A}$ of $M$ and define transition functions
@@
[…]
@@
[…]

### v2

@@ 8.1 Cauchy, Dirichlet and Neumann Boundary Conditions for the Laplace Equation
We denote by $\delta_{n}$  [^4] Not to be confused with the codifferential in the previous sections a sequence of regularizations of the Dirac delta function $\delta_{S}$ (it can be $S_{i}$, or $S_{o}$), such that $\lim_{n\to\infty}\delta_{n}=\delta_{S}$. Then, in the generalized Laplace equation (47), we choose (see Appendix B for justification):
@@
We observe that, in the singular limit $n\to\infty$, the regularization $\delta_{n}$ approaches the Dirac distribution $\delta_{S}$, supported on the neighborhood of the hypersurface $S$. In this limit, the terms involving derivatives of $\delta_{n}$ become highly localized near $S$, and act to impose both Dirichlet and Neumann conditions simultaneously on the surface.
As a result, equation (61) defines two decoupled Cauchy problems in $\mathbb{R}^{d}$: one on $\overline{\Omega_{i}}$ and one on $\overline{\Omega_{o}}$. Unless one solution is the harmonic continuation of the other, a discontinuity (in the solution or its gradient) appears at $S$. If only a Dirichlet condition $\phi_{S}=\phi_{d}$ is imposed, a discontinuity in the normal derivative $\nabla\phi_{S}\cdot\mathbf{n}$ may appear across $S$. Conversely, if only the Neumann condition $\nabla\phi_{S}\cdot\mathbf{n}=\nabla\phi_{d}\cdot\mathbf{n}$ is enforced, a jump in the solution may appear at $S$. These phenomena are closely related to classical potential theory, particularly the behavior of single-layer and double-layer potentials [21, 22].
Specifically, a single-layer potential yields a continuous potential across the interface but typically introduces a discontinuity in its normal derivative. In contrast, a double-layer potential maintains continuity of the normal derivative but allows for a jump in the potential itself [5]. The behavior of equation (61) in the limit $n\to\infty$ is thus analogous to a combined layer potential, where both Dirichlet and Neumann data are enforced.
${\mathbb{R}}^{d}$$\mathbf{\hat{n}}_{o}$$\mathbf{\hat{n}}_{i}$$\Omega_{i}$$\Omega_{o}$$S_{o}$$S_{i}$
[…]
Figure 4: Left: The closures $\overline{\Omega_{o}}$ and $\overline{\Omega_{o}}$ are considered as manifolds with boundary, sharing the boundary $S$. Right: A manifold with boundary is locally isomorphic (the isomorphism is denoted by $\psi$) with a half space ${\mathbb{H}}^{d}$.
Interestingly, equation (61) also provides a systematic framework for solving classical boundary value problems. Suppose one intends to solve a pure Dirichlet problem—either on the interior or exterior of $S$. In this case, one may first omit the gradient penalizing terms and consider the equation:
@@
Once $\phi_{n}$ is obtained from equation (62), the normal derivative $\nabla\phi_{n}\cdot\mathbf{n}$ on $S$ is obtained as well. In the context of elliptic boundary value problems, this is equivalent to applying the Dirichlet-to-Neumann (DtN) map $\Lambda$ (not to be confused with the homothetic transformation matrix in the previous sections), often denoted by
$\Lambda:\phi_{n}|_{S}\mapsto\nabla\phi_{n}\cdot\mathbf{n}|_{S}.$ (63)
Now, by assuming $\nabla\phi_{d}:=\nabla\phi_{n}$, the term $(\nabla\phi_{n}-\nabla\phi_{d})$ in the original equation (61) vanishes identically. Thus, one has in effect solved the complete equation (61) under a self-consistent assumption that the constructed $\phi_{d}$ is aligned with the solution gradient on $S$.
A similar argument holds for solving Neumann problems. In this case, one may omit the Dirichlet penalty term and retain only the gradient penalty term:
@@
This imposes a Neumann boundary condition by penalizing the difference between the gradient and the prescribed one $\nabla\phi_{d}$. Upon obtaining a solution $\phi_{n}$, one can define $\phi_{d}:=\phi_{n}$ on $S$, and again observe that the residual term $(\phi_{n}-\phi_{d})$ in equation (61) vanishes. Hence, the full equation is satisfied again under the assumption of gradient-data consistency.
This selective process regularizes the non-regular Cauchy problem and simplifies numerical treatment. It also clarifies the role of the full equation as a unified interface formulation, capable of encoding a range of classical boundary conditions. Moreover, the correspondence with the DtN operator can also be interpreted as a feature of the interface as a generalized boundary operator in the sense of nonlocal elliptic theory [6, 27].
8.2 A Non-Singular Model for a Point Electric Charge
@@
$E=\frac{1}{2}\int_{\mathbb{R}^{3}}|\nabla\phi|^{2}\,dV,$
diverges due to the integrable singularity at the origin. This divergence has been a foundational issue in classical field theory, prompting numerous attempts at regularization, ranging from distributional reformulations [12, 4] to the introduction of extended charge models [7, 2]. In the framework developed in the previous section, a new perspective emerges.
We consider a spherical interface $\partial B(R)\subset\mathbb{R}^{3}$, and solve the generalized Laplace equation with Dirichlet-type penalization:
$\nabla^{2}\phi_{n}+\frac{\Delta\delta_{n}}{1+\delta_{n}}(\phi_{n}-\phi_{d})=0,$ (62 revisited)
where $\delta_{n}$ is a regularization of the Dirac delta function supported on the surface of a ball, $\delta_{\partial B(R)}$, and $\phi_{d}$ is the prescribed Dirichlet potential
$\phi_{d}=\frac{C}{R},$
enforcing a Coulomb-like condition at $r=R$ .
Remark 2
It should be emphasized that the delta function is defined on the surface of the ball $\partial B(R)$, not at the origin. It makes an essentially different situation in comparison with a point delta function. A hollow sphere $S^{2}$ and a point are homotopically different objects irrespective of $R$, and they have different capacitance (from the potential theory viewpoint) irrespective of $R$.
Due to the highly localized nature of the penalty term, the solution $\phi_{n}$ adjusts sharply near the interface to match the boundary value, and decays as $\phi_{n}(r)\sim\frac{C}{r}$ for large $r$, consistent with the behavior of the classical potential in the far field.
@@
Figure 5: A non-singular point charge model with a spherical interface at $r=R$. The delta function is defined at the surface of the sphere. The interior potential is decoupled from the exterior, where the solution decays as $\phi_{n}\sim\frac{C}{r}$.
The structure of the equation ensures a natural decoupling of the interior and exterior domains. In the limit $n\to\infty$, the penalization enforces the boundary condition only on $\partial B(R)$, and the interior and exterior solutions become independent. Thus, the behavior of the solution in $\mathbb{R}^{3}\setminus B(R)$ is determined entirely by the enforced Dirichlet data and is unaffected by the field inside the ball.
Remark 3
It should be emphasized that we are not aware of the solution inside the ball $B(R)$. Inside the ball is another manifold in which the solution is determined from the boundary condition on its boundary.
We can now safely approach $R\to 0$, by which the ball $B(R)$ gets smaller, and the prescribed boundary potential $\phi_{d}=C/R$ grows, although neither $\partial B(R)$ changes to a point nor $\phi_{d}$ gets infinity. Moreover, since our domain is the exterior domain, and the potential inside the ball plays no role in the exterior solution, the energy,
$E=\frac{1}{2}\int_{\mathbb{R}^{3}\setminus B(R)}|\nabla\phi_{n}|^{2}\,dV,$ (65)
remains finite in the limit $R\to 0$, as the core is effectively excluded from the domain of integration.
Therefore, we have a physically faithful, Coulombian potential associated with a point-like source, but without the divergence of vacuum energy. The source is interpreted not as a singular point but as the limit of a penalized interface of vanishing size, carrying an enforced boundary potential. The infinite energy problem is thereby resolved through a combination of geometrically localized penalization and the decoupling of interior and exterior fields.
@@ 9 Conclusions
In this work, we have developed a Homothetic Gauge Theory (HGT) that systematically promotes local scale invariance to the status of a fundamental gauge principle, placing it on the same geometric footing as the interactions of the Standard Model. The central challenge of incorporating the non-linear operation of scaling into the linear framework of gauge theory was overcome by introducing a novel fiber-wise embedding into a doubled vector bundle. This construction linearizes the homothetic transformations, allowing them to be represented by a simple matrix group acting on the extended fibers.
Upon this foundation, we constructed the full machinery of a consistent gauge theory. A generalized Weyl-integrable geometry emerged, in which the spacetime metric is conformally rescaled by a dynamical scalar gauge field, $\lambda(x)$. The mathematical integrity of the theory was secured by developing a homothetic Hodge–de Rham complex, complete with generalized differential operators and a corresponding Hodge decomposition theorem. This framework ensures that the resulting field equations are well-posed and their solutions are unique under appropriate boundary conditions.
The primary physical application of HGT was demonstrated in the context of classical electromagnetism. The theory yields a set of homothetically covariant Maxwell equations that are inherently sourceless. Traditional source terms are replaced by geometric penalty terms that couple the electromagnetic field to the gradient of the scale field, $\nabla\lambda$. This unique feature allows boundary conditions to be encoded directly into the dynamics, reformulating boundary-value problems as sourceless field equations over the entire domain.
It has shown that local scale covariance is preserved by the HGT field equations while preferred scales arise from the selection of gauge/background data. In particular, the dilatation one–form $b=d\lambda$ and the offset fields $\phi_{d}$ act as order parameters: fixing $\lambda$ and $\phi_{d}$ selects solution sectors that break scale symmetry without introducing a Higgs-type potential. Thus, HGT realizes spontaneous symmetry breaking (SSB) as a property of solutions (and their data), not of the equations themselves.
Within homothetic electromagnetism (Secs. 56), the SSB viewpoint explains how physical scales enter the generalized Maxwell theory in a controlled way. The penalization framework and homothetic Lorenz gauge make the mechanism explicit: the penalization profile and the chosen $\lambda$ determine the emergent length scale, while the equations remain homothetically covariant. This solution-level SSB suggests several directions: coupling the mechanism to gravity (Cartan/Weyl-integrable interpretations), analyzing anomalies and quantization of the homothetic sector, studying dynamical selection of $\lambda$ and $\phi_{d}$ in matter-coupled models, and exploring additional interface limits beyond electrostatics. We expect these lines to sharpen the relation between mathematical scale covariance and the emergence of physical scales across HGT applications.
From the computational physicist’s viewpoint, perhaps the most significant physical consequence of this framework is a natural resolution to the classical problem of the infinite self-energy of a point charge. By modeling the source via a volume penalization method that enforces a Dirichlet condition in the limit, we constructed a non-singular electrostatic potential that is finite everywhere, including at the origin. The divergence is eliminated not by an ad-hoc modification of Maxwell’s laws, but as a direct consequence of the underlying geometry. This result suggests that what we perceive as sources may be geometric interfaces where the scale of spacetime undergoes rapid change.
Finally, by showing that the entire formalism can be understood as a specific gauge fixing within the broader context of conformal Cartan geometry, we have grounded HGT in a well-established mathematical tradition. The theory presented here is not merely a mathematical reformulation but a predictive framework with the potential for broad application. Future work could involve extending this formalism to non-Abelian gauge groups or exploring its implications for gravity, where the homothetic gauge field might offer new insights into dark energy or the cosmological constant problem.
@@ A.1 Conformal Cartan Geometry
  • the doubled bundle $TM\oplus TM$ of Section 2 is the direct sum of two tractor bundles carrying weights 0 and 1-1, respectively.
Hence all operators in $\S$ 2 are functorial in Cartan geometry.
@@ A.3 Interpretation of the Fixed Vector $X_{d}$
The canonical embedding $X\mapsto\bar{X}=\bigl{(}X~~X_{d}\bigr{)}^{\!\top}$ used in the main text corresponds, in the Cartan picture, to the choice of a weight 1-1 compensator tractor $X_{d}$. Its presence linearizes local scale transformations, exactly as a compensator field does in Weyl gravity.
@@ Appendix B Finding $\lambda$ for imposition of boundary conditions
Here, in the absence of a dynamical theory for $\lambda(x)$, we fix it in a way that allows us to apply boundary conditions. The result has the potential for applications in computational physics. The method presented here is not a complete solution to the mathematical problem of finding $\lambda$, but the $\lambda$ found is acceptable for sufficiently regular boundaries.
@@
$e^{\lambda}R\;=\;6\!\left[(\nabla\lambda)^{2}-\Box\lambda\right],\qquad e^{2\lambda}R_{\mu\nu}R^{\mu\nu}=\tfrac{3}{2}\!\bigl{[}(\nabla\lambda)^{2}-\Box\lambda\bigr{]}^{2},$
and analogous expressions for higher invariants. All of them remain finite because the square brackets are bounded. Furthermore, the affine–parameter integral along any causal geodesic, $\int\Omega^{2}\,ds,$ diverges only if $|\lambda|\to\infty$ faster than $\ln s$, which never happens for the present $\lambda$.
@@ Appendix C Principal $\mathcal{H}$-bundle formulation and curvature
Let $\mathcal{H}\cong({\mathbb{R}}_{>0},\cdot)$ be the homothetic group acting on each doubled fibre by block-diagonal multiplication. [^6] Locally—on a chart (U,x)(U,x)—a homothety is specified by a smooth function $\lambda\in C^{\infty}(U)$, the action being $v\mapsto e^{\lambda(x)}v$. Pointwise multiplication turns $\mathcal{H}$ into a $1$-dimensional abelian Lie group with Lie algebra $\mathfrak{h}\cong{\mathbb{R}}$ under addition. Choose a good open cover $\{U_{\alpha}\}_{\alpha\in A}$ of $M$ and define transition functions
@@
$\text{\emph{(i)}}\;\;\;R_{X^{\#}}\omega=1\quad\text{and}\quad\text{\emph{(ii)}}\;\;\;R_{h}^{*}\omega=Ad_{h^{-1}}\!\omega\;(=\omega\text{ since }\mathcal{H}\text{ is abelian}),$
@@
$R\;=\;d\Gamma+\Gamma\wedge\Gamma\quad(\text{Maurer–Cartan form}),$